OpenStudy (anonymous):

how wouldyou go about doing the equation 5y/3x*15x/25

OpenStudy (anonymous):

\[\frac{ 5y }{ 3x }\times \frac{ 15x }{ 25 }\] =\[\frac{ 75xy }{ 75x }\] Now if you simplify that, you get y. Also, please put this in the MATHEMATICS thread. You're posting in the wrong place.